# Grounding Multimodal LLMs to Embodied Agents that Ask for Help with Reinforcement Learning

### v2

@@ 3. Is it the <object_size> one?
Table 8: LLM prompt used for the fully observable text world graph with ReAct [53] zero-shot baseline.
1. nav(receptacle)
2. pick(object)
3. place(receptacle)
4. ask("question")
You can only ask the following 5 types of yes/no questions about the object in question:
1. Is it on the <receptacle>?
2. Is it the <object_instance>?
3. Is it the <object_size> one?
After you ask a question wait for the human to reply.
Here is a example of completing the task:
Instruction: Bring me bowl put it on cabinet
Receptacles with objects:
coffee table: blue casserole
dark table: red bowl, yellow bowl, black toy
light table: yellow dumbbell, blue bowl, red bowl
Observation action history:
Step 1:
Previous Actions: []
Agent at: Start
Thought: I see two red bowls, one yellow bowl and one blue bowl. It’s unclear which one user wants. I should ask about location of the bowl first and then the color to find the correct one.
Output Action: ask("Is it on the light table?")
Action Complete!
User response: no
Step 2:
Previous Actions: [ask("Is it on the light table?")]
Agent at: Start
Thought: User said it’s not on light table. This means user wants either a red or a yellow bowl. I should ask about the color of the bowl user wants.
Output Action: ask("Is it the red bowl?")
Action Complete!
User response: yes
Step 3:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?")]
Agent at: Start
Thought: User said it’s red bowl that is on the dark table. I will navigate to it.
Output Action: nav(dark table)
Action Complete!
Step 4:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table)]
Agent at: dark table
Thought: I need to pick the red bowl.
Output Action: pick(red bowl)
Action Complete!
Step 5:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl)]
Agent at: dark table
Thought: I need to navigate to cabinet to place it.
Output Action: nav(cabinet)
Action Complete!
Step 6:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet)]
Agent at: cabinet
Thought: I need to place red bowl on the cabinet.
Output Action: place(cabinet)
Action Complete!
Current observation:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet), place(cabinet)]
Agent at: cabinet
Thought: Task is done.
Output Action: Done()
Strictly follow the above format while solving given task.
Instruction: instruction
Receptacles: receptacles
Receptacles with objects:
receptacles_with_objects
Observation action history:
prev_observations
Current observation:
Previous Actions: prev_actions
Agent at: current_room
Thought:
Table 9: LLM prompt used for the fully observable text world graph with ReAct [53] baseline with in-context examples.
1. nav(receptacle)
2. pick(object)
3. place(receptacle)
4. ask("question")
You can only ask the following 5 types of yes/no questions about the object in question:
1. Is it on the <receptacle>?
2. Is it the <object_instance>?
3. Is it the <object_size> one?
After you ask a question wait for the human to reply.
Here is a example of completing the task:
Instruction: Bring me bowl put it on cabinet
Receptacles with objects:
coffee table: blue casserole
dark table: red bowl, yellow bowl, black toy
light table: yellow dumbbell, blue bowl, red bowl
Observation action history:
Step 1:
Previous Actions: []
Agent at: Start
Thought: I see two red bowls, one yellow bowl and one blue bowl. It’s unclear which one user wants. I should ask about location of the bowl first and then the color to find the correct one.
Output Action: ask("Is it on the light table?")
Action Complete!
User response: no
Step 2:
Previous Actions: [ask("Is it on the light table?")]
Agent at: Start
Thought: User said it’s not on light table. This means user wants either a red or a yellow bowl. I should ask about the color of the bowl user wants.
Output Action: ask("Is it the red bowl?")
Action Complete!
User response: yes
Step 3:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?")]
Agent at: Start
Thought: User said it’s red bowl that is on the dark table. I will navigate to it.
Output Action: nav(dark table)
Action Complete!
Step 4:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table)]
Agent at: dark table
Thought: I need to pick the red bowl.
Output Action: pick(red bowl)
Action Complete!
Step 5:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl)]
Agent at: dark table
Thought: I need to navigate to cabinet to place it.
Output Action: nav(cabinet)
Action Complete!
Step 6:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet)]
Agent at: cabinet
Thought: I need to place red bowl on the cabinet.
Output Action: place(cabinet)
Action Complete!
Current observation:
Previous Actions: [ask("Is it on the light table?"), ask("Is it the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet), place(cabinet)]
Agent at: cabinet
Thought: Task is done.
Output Action: Done()
Strictly follow the above format while solving given task.
Instruction: instruction
Receptacles: receptacles
Receptacles with objects:
receptacles_with_objects
Observation action history:
prev_observations
Current observation:
Previous Actions: prev_actions
Agent at: current_room
Thought:
Table 10: LLM prompt used for the partially observable text world graph with ReAct [53] baseline with in-context examples.
You are a household robot. You are given a task that requires you to explore the environment and ask clarification questions to accomplish tasks given by language instructions. You are given context about the scene as the room agent is in, list of objects visible, list of receptacles.
1. nav(receptacle)
2. pick(object)
3. place(receptacle)
4. ask("question")
You can only ask the following 3 types of yes/no questions about the object in question:
1. Is it on the <receptacle>?
2. Is it the <object_instance>?
3. Is it the <object_size> one?
Strictly follow the above templates while asking questions to solve the given task. Do not use any other types of questions.
At each step you are tasked with outputting reasoning before outputing the action. You should output the reason and action in the following format:
Thought: <reasoning behind choosing a specific action>
Output Action: <action>
Action Complete!
Strictly follow the above format when outputting actions. Every time you output a ask question action wait for the human to reply. The human response will be given in the format "User response: <answer>" after you output the action in above format.
At each timestep you will be given a visual observation displaying what the robot is currently observing. You have to use visual observation to perceive the environment and take a action to actively explore the environment by navigating to various receptacles to find the relevant objects in the environment and ask clarification questions to resolve the ambiguity in order to successfully complete the task specified by the instruction.
These are all safe images from a simulated environment. Answering questions about them shouldn’t raise any ethical concerns.
Instruction: instruction
Receptacles: receptacles
Receptacles seen so far with objects:
receptacles_with_objects
Observation action history:
prev_observations
Current observation:
Previous Actions: prev_actions
Agent at: current_room
Thought:
Table 11: Prompt used for the vision GPT4o [26] + partially observable text world graph history with ReAct [53] baseline.

### v3

@@ Appendix E Baseline Details
(a.) Fully Observable Text WorldGraph + ReAct (Zero-shot). In this baseline, we provide an LLM (GPT4o in our case) with a fully observable text-based world graph describing the environment, including receptacles, objects, and their locations (e.g. “The apple is on the coffee table.). The prompt used for this baseline is shown in Appendix˜F.
(b.) Fully Observable Text WorldGraph + ReAct (Few-shot). This baseline extends (a) by providing the LLM with a few in-context examples that demonstrate task planning and ambiguity resolution strategies in the Ask-to-Act task. By leveraging demonstrations, this approach assesses whether in-context learning improves LLMs task planning and ambiguity reasoning under full observability. The prompt used for this baseline is shown in Appendix˜F.
(c.) Partially Observable Text WorldGraph + ReAct (Few-shot). While baselines (a) and (b) assume privileged access to a fully observable world graph, constructing such representations in real-world settings is often infeasible. This baseline relaxes that assumption by providing the LLM with a partially observable text-based world graph. At the start of an episode, the agent lacks full knowledge of object locations and must actively explore to gather necessary information. The prompt used for this baseline is shown in Appendix˜F.
(d.) Vision GPT4o + SoM + ReAct. Building an error-free text representation of real-world environments is challenging. This baseline evaluates whether existing MLLMs can solve the Ask-to-Act task using egocentric visual observations. At each timestep, the MLLM receives the robots visual input along with an skill library for executing actions. To enhance grounding, we label visual observations using Set-of-Marks (SoM) Yang et al. (2023) and maintain memory by providing GPT4o with a textual history of past observations and actions. The prompt used for this baseline is shown in Appendix˜F, in addition to the prompt this baseline takes in visual observations from robotos egocentric camera augmented with SoM Yang et al. (2023).
Appendix F Limitations
In this work, the AutoAsk method is purposefully constrained to questions relevant to the Ask-to-Act task to simplify the learning problem for a small scale 0.50.5B model. We find that using completely unconstrained question sampling (i.e. without explicitly including useful question templates in the prompt) leads to significantly poor performance for all finetuned and zero-shot methods. We believe this limitation can be addressed by using larger scale models (3B, 7B or 72B) which can be explored as part of future work. Another limitation of our work is the limited evaluation setup of $7$ tasks used in Ask-to-Act which is not a complete representative for real-world scenarios. An important future work is to study a more realistic setup of Ask-to-Act with a much larger and more complex set of ambiguous tasks and an open-ended question-answer setup. There are a range of open research questions required to do this, including the use (and potentially training/distillation) of efficient LLM-as-reward models that can be used within an RL pipeline. Further study is also needed to determine exactly how accurate or reliable the rewards must be to train effective models. As it is anticipated that LLM-generated rewards will not be as accurate in less constrained settings. These aspects are limitations of our current work but exciting research directions for future work.
For each task given as a language instruction you have to output a sequence of actions that the agent should take. These actions can include a clarification question as well. Ask a question only when required. For each action also output the reason to take the action with it. Actions can be one of the following: 1. nav(receptacle) 2. pick(object) 3. place(receptacle) 4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it 5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it 6. ask(question) You can ask the following 9 types of questions about the object in question: 1. Is target object on the <receptacle>? 2. Is <object_instance> the target object? 3. Is target object the <object_size> one? 4. Where is the <object_category> located? 5. What color is <object_category>? 6. Can you describe the <object_category>? 7. Is <object_instance> clutter? 8. Are <object_category> clutter? 9. Which <receptacle> to place the <object_instance> on/in? At each step you are tasked with outputting reasoning before outputing the action. You should output the reason and action in the following format: Thought: <reasoning behind choosing a specific action> Output Action: <action> Action Complete! Strictly follow the above format when outputting actions. Every time you output a ask question action wait for the human to reply. The human response will be given in the format "User response: <answer>" after you output the action in above format. Strictly follow the above format while solving given task. Instruction: {instruction} Receptacles: {receptacles} Receptacles with objects: {receptacles_with_objects} Observation action history: {prev_observations} Current observation: Previous Actions: {prev_actions} Agent at: {current_room} Thought:
For each task given as a language instruction you have to output a sequence of actions that the agent should take. These actions can include a clarification question as well. Ask a question only when required. For each action also output the reason to take the action with it. Actions can be one of the following: 1. nav(receptacle) 2. pick(object) 3. place(receptacle) 4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it 5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it 6. ask(question) You can ask the following 9 types of questions about the object in question: 1. Is target object on the <receptacle>? 2. Is <object_instance> the target object? 3. Is target object the <object_size> one? 4. Where is the <object_category> located? 5. What color is <object_category>? 6. Can you describe the <object_category>? 7. Is <object_instance> clutter? 8. Are <object_category> clutter? 9. Which <receptacle> to place the <object_instance> on/in? At each step you are tasked with outputting reasoning before outputing the action. You should output the reason and action in the following format: Thought: <reasoning behind choosing a specific action> Output Action: <action> Action Complete! Strictly follow the above format when outputting actions. Every time you output a ask question action wait for the human to reply. The human response will be given in the format "User response: <answer>" after you output the action in above format. After you ask a question wait for the human to reply. Here are two examples of completing the task: Example 1: Instruction: Bring me bowl put it on cabinet Receptacles: [light table, chair, sofa, dark table, tv stand, cabinet, sink, fridge, right counter drawer, left counter drawer] Receptacles with objects: coffee table: blue casserole dark table: red bowl, yellow bowl, black toy light table: yellow dumbbell, blue bowl, red bowl Observation action history: Step 1: Previous Actions: [] Agent at: Start Thought: I see two red bowls, one yellow bowl and one blue bowl. Its unclear which one user wants. I should ask about location of the bowl first and then the color to find the correct one. Output Action: ask("Is the target object on the light table?") Action Complete! User response: no Step 2: Previous Actions: [ask("Is the target object on the light table??")] Agent at: Start Thought: User said its not on light table. This means user wants either a red or a yellow bowl. I should ask about the color of the bowl user wants. Output Action: ask("Is the target object the red bowl?") Action Complete! User response: yes Step 3: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?")] Agent at: Start Thought: User said its red bowl that is on the dark table. I will navigate to it. Output Action: nav(dark table) Action Complete! Step 4: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?"), nav(dark table)] Agent at: dark table Thought: I need to pick the red bowl. Output Action: pick(red bowl) Action Complete! Step 5: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?"), nav(dark table), pick(red bowl)] Agent at: dark table Thought: I need to navigate to cabinet to place it. Output Action: nav(cabinet) Action Complete! Step 6: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet)] Agent at: cabinet Thought: I need to place red bowl on the cabinet. Output Action: place(cabinet) Action Complete! Current observation: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet), place(cabinet)] Agent at: cabinet Thought: Task is done. Output Action: Done() Example 2: Instruction: Clear clutter from the sofa and put it on the cabinet Receptacles: [light table, chair, sofa, dark table, tv stand, cabinet, sink, fridge, right counter drawer, left counter drawer] Receptacles with objects: coffee table: blue casserole sofa: red bowl, black toy, blue casserole light table: yellow dumbbell, blue toy Observation action history: Step 1: Previous Actions: [] Agent at: Start Thought: I dont know what the user means by clutter. I should first identify which objects on couch are user thinks are clutter. Output Action: ask("Is red bowl clutter?") Action Complete! User response: yes Step 2: Previous Actions: [ask("Is red bowl clutter?")] Agent at: Start Thought: User said red bowl is clutter. There could be more objects which are clutter so I should ask about it. Output Action: ask("Is blue casserole clutter?") Action Complete! User response: yes Step 3: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?")] Agent at: Start Thought: User said red bowl and blue casserole are clutter. There could be more objects which are clutter so I should ask about it. Output Action: ask("Is black toy clutter?") Action Complete! User response: no Step 4: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), ask("Is black toy clutter?")] Agent at: Start Thought: Looks like only red bowl and blue casserole are clutter and need to be moved to cabinet. Output Action: nav(sofa) Action Complete! Step 5: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa)] Agent at: sofa Thought: I need to pick the red bowl and then blue casserole. Output Action: pick(red bowl) Action Complete! Step 6: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl)] Agent at: cabinet Thought: I need to navigate to cabinet to place red bowl on the cabinet. Output Action: nav(cabinet) Action Complete! Step 7: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl), nav(cabinet)] Agent at: cabinet Thought: I need to place red bowl on the cabinet. Output Action: place(cabinet) Action Complete! Step 8: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl), nav(cabinet), place(cabinet)] Agent at: cabinet Thought: I need to go back to sofa to pick blue casserole. Output Action: nav(sofa) Action Complete! Step 9: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl), nav(cabinet), place(cabinet), nav(sofa)] Agent at: sofa Thought: I need to pick blue casserole. Output Action: pick(blue casserole) Action Complete! Step 10: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl), nav(cabinet), place(cabinet), nav(sofa), pick(blue casserole)] Agent at: sofa Thought: I need to navigate to cabinet to place blue casserole on the cabinet. Output Action: nav(cabinet) Action Complete! Current observation: Previous Actions: [ask("Is red bowl clutter?"), ask("Is blue casserole clutter?"), nav(sofa), pick(red bowl), nav(cabinet), place(cabinet), nav(sofa), pick(blue casserole), nav(cabinet)] Agent at: cabinet Thought: I need to place blue casserole on the cabinet. Output Action: place(cabinet) Action Complete! Current observation: Previous Actions: [ask("Is the target object on the light table??"), ask("Is the target object the red bowl?"), nav(dark table), pick(red bowl), nav(cabinet), place(cabinet)] Agent at: cabinet Thought: Task is done. Output Action: Done() Strictly follow the above format while solving given task. Instruction: {instruction} Receptacles: {receptacles} Receptacles with objects: {receptacles_with_objects} Observation action history: {prev_observations} Current observation: Previous Actions: {prev_actions} Agent at: {current_room} Thought:
For each task given as a language instruction you have to output a sequence of actions that the agent should take. These actions can include a clarification question as well. Ask a question only when required. For each action also output the reason to take the action with it. Actions can be one of the following: 1. nav(receptacle) 2. pick(object) 3. place(receptacle) 4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it 5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it 6. ask(question) You can ask the following 9 types of questions about the object in question: 1. Is target object on the <receptacle>? 2. Is <object_instance> the target object? 3. Is target object the <object_size> one? 4. Where is the <object_category> located? 5. What color is <object_category>? 6. Can you describe the <object_category>? 7. Is <object_instance> clutter? 8. Are <object_category> clutter? 9. Which <receptacle> to place the <object_instance> on/in? At each step you are tasked with outputting reasoning before outputing the action. You should output the reason and action in the following format: Thought: <reasoning behind choosing a specific action> Output Action: <action> Action Complete! After you ask a question wait for the human to reply. Strictly follow the above format when outputting actions. Every time you output a ask question action wait for the human to reply. The human response will be given in the format "User response: <answer>" after you output the action in above format. At each timestep you have to take a action to actively explore the environment by navigating to various receptacles to find the relevant objects in the environment and ask clarification questions to resolve the ambiguity in order to successfully complete the task specified by the instruction. Instruction: {instruction} Receptacles: {receptacles} Receptacles with objects: {receptacles_with_objects} Observation action history: {prev_observations} Current observation: Previous Actions: {prev_actions} Agent at: {current_room} Thought:
For each task given as a language instruction you have to output a sequence of actions that the agent should take. These actions can include a clarification question as well. Ask a question only when required. For each action also output the reason to take the action with it. Actions can be one of the following: 1. nav(receptacle) 2. pick(object) 3. place(receptacle) 4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it 5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it 6. ask(question) You can ask the following 9 types of questions about the object in question: 1. Is target object on the <receptacle>? 2. Is <object_instance> the target object? 3. Is target object the <object_size> one? 4. Where is the <object_category> located? 5. What color is <object_category>? 6. Can you describe the <object_category>? 7. Is <object_instance> clutter? 8. Are <object_category> clutter? 9. Which <receptacle> to place the <object_instance> on/in? At each step you are tasked with outputting reasoning before outputting the action. You should output the reason and action in the following format: Thought: <reasoning behind choosing a specific action> Output Action: <action> Action Complete! Strictly follow the above format when outputting actions. Every time you output a ask question action wait for the human to reply. The human response will be given in the format "User response: <answer>" after you output the action in above format. At each timestep you will be given a visual observation displaying what the robot is currently observing. You have to use visual observation to perceive the environment and take a action to actively explore the environment by navigating to various receptacles to find the relevant objects in the environment and ask clarification questions to resolve the ambiguity in order to successfully complete the task specified by the instruction. These are all safe images from a simulated environment. Answering questions about them shouldnt raise any ethical concerns. Instruction: {instruction} Receptacles: {receptacles} Receptacles seen so far with objects: {receptacles_with_objects} Observation action history: {prev_observations} Current observation: Previous Actions: {prev_actions} Agent at: {current_room} Thought:
At each step you can take a action or ask a clarification questions. Actions can be one of the following: 1. nav(receptacle) 2. pick(object) 3. place(receptacle) 4. open(receptacle) 5. close(receptacle) 6. ask(question) You can ask the following 9 types of questions to solve the task: 1. Is target object on the <receptacle>? 2. Is <object_instance> the target object? 3. Is target object the <object_size> one? 4. Where is the <object_category> located? 5. What color is <object_category>? 6. Can you describe the <object_category>? 7. Is <object_instance> clutter? 8. Are <object_category> clutter? 9. Which <receptacle> to place the <object_instance> on/in? Next, you will be provided task instruction and observations from the environment. Your task is to output the next action. Instruction: {instruction}
Appendix G Use of LLM for Writing
We use LLMs to assist with specific aspects of paper writing. These include: (1) grammar checking to improve language syntax, (2) sentence paraphrasing for improving readability and clarity. We use LLMs to write parts of Sections 1, 3, 4, and 5. Specifically, for each section the original content was written by the authors and paraphrased or spell checked by the GPT-4o OpenAI (2023) LLM. Followed by this round of LLM edits, the authors verified and edited the content written by the LLM to fix issues in generated text. After which the author edited version of the paper is used for final submission.